# Safe On-Orbit Dislodging of Deployable Structures via Robust Adaptive MPC

### v1

@@ III-A Linear Parametric Modeling
To facilitate learning the uncertainty parameters $\boldsymbol{\rho}_{k}$, we reorganize the hinge dynamics (1) into the linear parametric state-space
@@
where $\mathbf{I}\in\mathbb{R}^{2\times 2}$ is the identity matrix, $\mathbf{x}_{k}=[\theta_{k},\dot{\theta}_{k}]^{\top}{\in}\mathbb{R}^{2}$ and $\mathbf{w}_{k}=\left[0,\tfrac{\tau_{s_{k}}}{M}\right]^{\top}{\in}\mathbb{R}^{2}$ denotes the disturbance which is bounded. $h\in(0,1)$ is the sampling period for discretization of the state matrices $\mathbf{A}(\boldsymbol{\rho}_{k})$ and $\mathbf{B}(\boldsymbol{\rho}_{k})$, in which can be parameterized as
@@
Then based on (1), we can get (7) in detail as
@@
Note that considering the gripper may not be perpendicular to the rod for all time during the dislodging task, $\alpha$ is also time-varying. Also, we assume the $\alpha$ would not deviate more than $70^{\circ}$ from the initial position and based on $\sin\theta=\sin(\pi+\phi)\approx-\phi$ and $\cos\theta=\cos(\pi+\phi)\approx-1$, we can use $-\phi$ and $-1$ to replace $\sin(\alpha_{k})$ and $\cos(\alpha_{k})$ respectively to get $\mathcal{D}_{k}=-d_{g}\cdot\phi_{k}$ and $\mathcal{L}_{k}=-1$.
@@ III-B Set-Based Parameter Estimation
In this section, we present our set-based parameter estimator for dislodging in which bounds the uncertain parameters $\boldsymbol{\rho}_{k}$ of the linear parametric model (6) using real-time state $\mathbf{x}_{k}$ of the client and control input $\mathbf{u}_{k}$ from the servicer. Our robust adaptive MPC will leverage these bounds to ensure robust constraint satisfaction despite parametric uncertainty. To bound the parameter uncertainty, we make the following assumption about the boundedness of the measurement noise $\mathbf{w}_{k}$.
@@
The disturbance set $\mathbb{W}$ is a bounded polytope described by the $n_{w}$ constraints in the set
@@
where $\mathbf{H}_{w}\in\mathbb{R}^{n_{w}\times 2}$ and $\boldsymbol{\delta}_{w}\in\mathbb{R}^{n_{w}}$ The parameter set $\boldsymbol{\rho}_{k}$ are iteratively updated at each time step $k$ to form a set $\mathcal{P}_{k}$ which bounds all possible values of the uncertain parameters $\boldsymbol{\rho}_{k}$. This update is achieved by constructing a non-falsified parameter set, utilizing measurement data from the preceding $s$ time steps $\nu_{s}$, as follows:
$\Delta_{k}$ :=$:=\biggr{\{}\boldsymbol{\rho}_{k}\in\mathbb{R}^{p}\>\biggr{|}\mathbf{x}_{\nu_{s}+1}-\left(\mathbf{A}(\boldsymbol{\rho}_{k})\mathbf{x}_{\nu_{s}}+\mathbf{B}(\boldsymbol{\rho}_{k})\mathbf{u}_{\nu_{s}}\right)\in\mathbb{W}\biggr{\}}$ (9)
=$=\biggr{\{}\boldsymbol{\rho}_{k}\in\mathbb{R}^{p}\>\biggr{|}\>-\mathbf{H}_{w}\mathbf{D}_{{\nu}_{s}{-}{1}}\boldsymbol{\rho}_{k}\leq\boldsymbol{\delta}_{w}+\mathbf{H}_{w}\mathbf{d}_{k}\biggr{\}},$
where $\forall\nu_{s}{\in}\{k{-}s,k{-}1\}_{\mathbb{N}}$ denotes the preceding time steps from time $k{-}s$ to $k{-}1$, and $\mathbf{D}_{\nu_{s}}{\in}\mathbb{R}^{n\times 3}$ can be defined as
@@
In (9), the non-falsified set $\Delta_{k}$ defines the set of all parameters that could have generated the measurement sequence $\{\mathbf{x}_{k-s},\ldots,\mathbf{x}_{k}\}$. To manage computational complexity, the polytopic set $\mathcal{P}_{k}$ is defined using a fixed number of linear constraints
@@
where the fixed matrix $\mathbf{H}_{\rho_{k}}\in\mathbb{R}^{n_{p}\times 3}$ is chosen offline and $\boldsymbol{\delta}_{\rho_{k}}\in\mathbb{R}^{n_{p}}$ is updated online. To account for the time-varying (3) parameters $\boldsymbol{\rho}_{k}^{*}$, we introduce a dilation operator with $\varsigma\in\mathbb{N}$ for $\mathcal{P}$ as
@@
at time step $k$. Our robust adaptive MPC requires that we predict how the bounds $\hat{\mathcal{P}}_{\varsigma|k}$ on the time-varying parameters $\boldsymbol{\rho}_{k}$ evolve over the prediction horizon. These prediction bounds are given by
@@
for $\varsigma=1,\dots,N-1$ where $\hat{\mathcal{P}}_{\varsigma\mid k}$ is the predicted uncertainty bound at time $k+\varsigma$.
@@
$\begin{split}[\boldsymbol{\delta}_{\rho_{k}}]_{i}\>&=\>\min_{\boldsymbol{\rho}\in\mathbb{R}^{p}}\quad[\boldsymbol{\delta}_{\rho}]_{i}\boldsymbol{\gamma}_{i},\quad\forall i\in\{1,n_{\rho^{+}}\}_{\mathbb{N}}\\ &\text{s. t. }\quad\begin{bmatrix}\mathbf{H}_{\rho}\\ -\mathbf{H}_{w}\mathbf{D}_{\nu_{s}}\end{bmatrix}\boldsymbol{\rho}^{+}\leq\begin{bmatrix}\boldsymbol{\delta}_{\rho_{k-1}}\\ \boldsymbol{\gamma}+\mathbf{H}_{w}\boldsymbol{d}_{k+1}\end{bmatrix}.\end{split}$ (15)
@@ III-C Robust Tube Constraints
In this section, we derive a state-space tube $\mathbb{X}_{\varsigma|k}$ for $\varsigma=1,\dots,N$ where $N$ is the prediction horizon and we can guarantee contains the state $\mathbf{x}_{k}$ despite parametric uncertainty $\boldsymbol{\rho}_{k}\in\mathcal{P}_{k}$ and disturbances $\mathbf{w}_{k}\in\mathbb{W}$. By remaining in the state tube, we can guarantee that the servicer would not damage the client. The states of the client and the control inputs from the servicer must satisfy the constraints as
$\mathbb{Z}$ =$=\left\{(\mathbf{x}_{k},\mathbf{u}_{k})\in\mathbb{R}^{n}\times\mathbb{R}^{m}\bigr{|}\mathbf{F}\mathbf{x}_{k}+\mathbf{G}\mathbf{u}_{k}\leq\boldsymbol{\gamma}\right\},$ (16)
where the matrices $\mathbf{F}\in\mathbb{R}^{n_{\gamma}\times n}$ and $\mathbf{G}\in\mathbb{R}^{n_{\gamma}\times n}$ are derived from the input (4a) and state (4b) constraints.
The tube MPC approach proposed in [29] ensures robust constraint satisfaction. The control input is parameterized using a feedback gain $\mathbf{K}\in\mathbb{R}^{m\times n}$ as
@@
A gain $\mathbf{K}$ satisfying Assumption 2 can be computed using standard robust control techniques, e.g., following the approach in [30].
@@
for $j=0,\dots,N-1$ and for all parameters $\boldsymbol{\rho}_{k}\in\mathcal{P}_{\varsigma|k}$. This ensures that $\mathbf{x}_{\varsigma|k}\in\mathbb{X}_{\varsigma|k}$ for all the realizations of uncertainty and disturbance. To manage computational complexity, the tube cross-section at each time step, $\left\{\mathbb{X}_{\varsigma\mid k}\right\}_{j\in\{0,N\}_{\mathbb{N}}}$, is parameterized by translating and scaling of the set
@@
$\mathbb{X}_{\varsigma|k}&=\{\mathbf{z}_{\varsigma|k}\}\oplus\mathbb{X}_{0,k}^{\curlyvee}=\{\mathbf{x}\in\mathbb{R}^{n}|\mathbf{H}_{x,k}^{\curlyvee}(\mathbf{x}-\mathbf{z}_{\varsigma|k})\leq\boldsymbol{\gamma}\}\\ &=\{\mathbf{z}_{\varsigma|k}\}\oplus\vartheta_{\varsigma|k}\text{co}\{\mathbf{x}^{1},\mathbf{x}^{2},\ldots,\mathbf{x}^{v}\}.\end{aligned}$
@@ III-D Reformulation of Robust Tube Constraints
From [31], we can reformulate the robust tube constraints (20) as linear equality and inequality constraints as follows. Let the state tube $\{\mathbb{X}_{\varsigma|k}\}_{\varsigma\in\{{0},{N}\}_{\mathbb{N}}}$ be parameterized according to (20). Then, the constraints (16) and set-dynamics (18) are satisfied if and only if $\forall j\in\{1,v\}_{\mathbb{N}}$, $\varsigma\in\{0,N{-}1\}_{\mathbb{N}}$ there exists $\boldsymbol{\Gamma}_{\varsigma|k}^{j}\in\mathbb{R}^{n_{x}\times n_{p}}_{+}$ such that
@@ III-E Robust Terminal Set
[…]
@@
[…]
@@
Assumption 3 implies that the set $\mathbb{X}_{T}$ is a robust positively invariant (RPI) set for the set-dynamics in $(\mathbf{z},\vartheta)$, with an additional constraint that the set $\mathbb{X}_{N|k}$ remains centered at origin. Note that Assumption 2 is a necessary condition for Assumption 3 to be satisfied, but they are stated separately to emphasize that the stronger assumption is only needed to implement the terminal condition.
@@ III-F Robust Adaptive MPC Algorithm
with =$\boldsymbol{\mho}=\bigl{\{}\{\mathbf{z}_{\varsigma|k}\},\{\vartheta_{\varsigma|k}\},\{\mathbf{v}_{\varsigma|k}\},\{\mathbf{K}^{j}_{\varsigma|k}\}\bigl{\}}$ as the decision variable. $\ell\left(\mathbb{X}_{i\mid k},\mathbf{v}_{i\mid k}\right)=\max_{\mathbf{x}\in\mathbb{X}}||\mathbf{Q}\mathbf{x}||_{\infty}+||\mathbf{R}(\mathbf{K}\mathbf{x}+\mathbf{v})||_{\infty}+||\boldsymbol{\Upsilon}\boldsymbol{\rho}||_{\infty}$, and $\mathbf{Q},\mathbf{R},\boldsymbol{\Upsilon}$ are positive definite matrices. Note that in our work, we introduce the term $||\boldsymbol{\Upsilon}\boldsymbol{\rho}||_{\infty}$ which explicitly optimizing not only for the system’s control performance but also for improving the accuracy of your time-varying parameter estimates. The results for parameter estimation in Section IV evidently shows the advantage of this novel contribution compared with traditional adaptive MPC. A linear cost function is selected to enable its reformulation through linear inequalities resulting in the MPC optimization problem becoming a linear program.
@@
Theorem 1: Let the assumptions 1, 2 and 3 be satisfied and an initial feasible solution exist. Let the parameter estimate set $\mathcal{P}_{k}$ be defined by the set dynamics (13) and admissible input set $\mathcal{I}_{v_{s}}\left(\mathbf{x}_{k},\mathcal{P}_{k}\right)$ as
@@
Then, the closed loop system using Algorithm 1 satisfies the following properties for all $k>k_{0}$:
@@
For Property (1), we know that $\boldsymbol{\rho}_{k}\in\mathcal{P}_{k}$ and define $\epsilon_{\rho_{k}}=\boldsymbol{\rho}_{k+1}-\boldsymbol{\rho}_{k}$. Then we can get
@@
$\tilde{\mathbf{H}}_{\rho_{k}}\boldsymbol{\rho}_{k+1}=\tilde{\mathbf{H}}_{\rho_{k}}\boldsymbol{\rho}_{k}+\tilde{\mathbf{H}}_{\rho_{k}}\epsilon_{\rho_{k}}\leq\tilde{\boldsymbol{\delta}}_{\rho_{k}}+d_{\rho}\boldsymbol{\gamma}.$
Since by assumption $\boldsymbol{\rho}_{k+1}\in\mathcal{P}$, we can get $\boldsymbol{\rho}_{k+1}\in\mathbf{d}_{\varsigma}\left(\mathcal{P}_{k}\cap\Delta_{k+1}\right)$.
@@
$\tilde{\mathbb{X}}_{\varsigma+1\mid k+1}\supseteq\mathbf{A}_{\text{cl}}(\boldsymbol{\rho})\tilde{\mathbb{X}}_{\varsigma\mid k+1}\oplus\mathbf{B}(\theta)\tilde{\mathbf{v}}_{\varsigma\mid k+1}\oplus\mathbb{W}.$
where $\tilde{\mathbf{v}}_{\varsigma\mid k+1}=\mathbf{v}_{\varsigma+1\mid k}^{*}$ and $\varsigma\in\{0,N{-}2\}_{\mathbb{N}}$. Based on Assumption 3 that there exist $\mathbb{X}_{N\mid k+1}$ satisfy the last set when $\varsigma=N{-}1$, that proof that $\mathcal{I}_{v_{s}}\neq\{\varnothing\}$ during all time step.
Property (3) is the direct result of Section III-D.

### v2

@@ III-A Linear Parametric Modeling
To facilitate learning the uncertainty parameters $\boldsymbol{\rho}_{k}$ and build up the connection for the solar panel between the rotation along piovt axis relative to frame $\mathcal{L}$ in 1 DoF and the motion relative to the frame $\mathcal{S}$ in 6 DoF, we reorganize the hinge dynamics (1) into the linear parametric state-space
@@
where $\mathbf{I}\in\mathbb{R}^{2\times 2}$ is the identity matrix, $\mathbf{x}_{k}=[\theta_{k},\dot{\theta}_{k}]^{\top}{\in}\mathbb{R}^{2}$ and $\mathbf{w}_{k}=\left[0,\tfrac{\tau_{m_{k}}}{M}\right]^{\top}{\in}\mathbb{R}^{2}$ denotes the disturbance which is bounded. $h\in(0,1)$ is the sampling period for discretization of the state matrices $\mathbf{A}(\boldsymbol{\rho}_{k})$ and $\mathbf{B}(\boldsymbol{\rho}_{k})$, which can be parameterized as
@@
where $n=3$ denotes the number of unvertain parameter in $\boldsymbol{\rho}_{k}$. Then based on (1), we can get (7) in detail as
@@
Note that considering the gripper may not be perpendicular to the rod for all time during the dislodging task, $\alpha$ is also time-varying. Here, we choose $\phi$ to denote the deviation of the EE’s angle from its ideal position which is perpendicular to the central axis of the solar panel. Also, we assume the $\alpha$ would not deviate more than $70^{\circ}$ from the initial position so that $\phi=90^{\circ}-\alpha\leq 20^{\circ}$. Under this assumption, based on $\sin\theta=\sin(\pi+\phi)\approx-\phi$ and $\cos\theta=\cos(\pi+\phi)\approx-1$, we can use $\phi$ and 1-1 to replace $\sin(\alpha_{k})$ and $\cos(\alpha_{k})$ respectively to get $\mathcal{D}_{k}=-d_{g}\cdot\phi_{k}$ and $\mathcal{L}_{k}=-1$.
@@ III-B Set-Based Parameter Estimation
In this section, we present our set-based parameter estimator for dislodging which bounds the uncertain parameters $\boldsymbol{\rho}_{k}$ of the linear parametric model (6) using real-time state $\mathbf{x}_{k}$ of the client and control input $\mathbf{u}_{k}$ from the servicer. Our robust adaptive MPC will leverage these bounds to ensure robust constraint satisfaction despite parametric uncertainty. To bound the parameter uncertainty, we make the following assumption about the boundedness of the measurement noise $\mathbf{w}_{k}$.
@@
The disturbance set $\mathbb{W}$ is a bounded polytope which includes exogenous disturbances and small unmodeled actuator nonlinearities, e.g., deadzones, backlash, etc, at the end-effector or hybrid hinge as described by the $n_{w}$ constraints in the set
@@
where $\mathbf{H}_{w}\in\mathbb{R}^{n_{w}\times 2}$ and $\boldsymbol{\delta}_{w}\in\mathbb{R}^{n_{w}}$. Note that we assume that no deadzone nonlinearity is modeled explicitly, and its effect is treated as an additive bounded disturbance within $\mathbb{W}$. The parameter set $\boldsymbol{\rho}_{k}$ is iteratively updated at each time step $k$ to form a set $\mathcal{P}_{k}$ which bounds all possible values of the uncertain parameters $\boldsymbol{\rho}_{k}$. This update is achieved by constructing a non-falsified parameter set, utilizing measurement data from the preceding $s$ time steps $\nu_{s}$, as follows:
$\Delta_{k}$ $=\biggr\{\boldsymbol{\rho}_{k}\in\mathbb{R}^{p}\>\biggr|\mathbf{x}_{\nu_{s}+1}-\left(\mathbf{A}(\boldsymbol{\rho}_{k})\mathbf{x}_{\nu_{s}}+\mathbf{B}(\boldsymbol{\rho}_{k})\mathbf{u}_{\nu_{s}}\right)\in\mathbb{W}\biggr\}$ (9)
$=\biggr\{\boldsymbol{\rho}_{k}\in\mathbb{R}^{p}\>\biggr|\>-\mathbf{H}_{w}\mathbf{D}_{{\nu}_{s}{-}{1}}\boldsymbol{\rho}_{k}\leq\boldsymbol{\delta}_{w}+\mathbf{H}_{w}\mathbf{d}_{k}\biggr\},$
where $\forall\nu_{s}{\in}\{k{-}s,k{-}1\}_{\mathbb{N}}$ denotes the preceding time steps from time ksk{-}s to k1k{-}1, and $\mathbf{D}_{\nu_{s}}{\in}\mathbb{R}^{n\times 3}$ can be defined as
@@
In (9), the non-falsified set $\Delta_{k}$ is constructed by performing a consistency check for every candidate parameter against the latest measurement data. This is achieved by using the measured sequence $\{\mathbf{x}_{k-s},\ldots,\mathbf{x}_{k}\}$ and inputs to calculate the disturbance $\mathbf{w}_{k}$ that a candidate parameter $\boldsymbol{\rho}_{k}$ would imply. The parameter is then ’falsified’ if this disturbance falls outside the known bounded set $\mathbb{W}$. This process results in a set of linear inequalities that geometrically define $\Delta_{k}$ as a computationally tractable polytope, representing all parameters consistent with the observed system behavior. To manage computational complexity, the polytopic set $\mathcal{P}_{k}$ is defined using a fixed number of linear constraints
@@
where the fixed matrix $\mathbf{H}_{\rho_{k}}\in\mathbb{R}^{n_{p}\times 3}$ is chosen offline and $\boldsymbol{\delta}_{\rho_{k}}\in\mathbb{R}^{n_{p}}$ is updated online. Note that as new measurements become available, the number of constraints defining the parameter set would grow indefinitely, making the problem computationally intractable for a real-time controller. To manage this computational complexity, the polytopic set $\mathcal{P}_{k}$ in (11) using a fixed number of linear constraints, where the matrix $\mathbf{H}_{\rho_{k}}$ chosen offline. Rather than adding new constraints with each measurement, the algorithm simply updates the vector $\boldsymbol{\delta}_{\rho_{k}}$ online, which tightens this fixed-structure polytope based on new information. To account for the time-varying (3) parameters $\boldsymbol{\rho}_{k}^{*}$, we introduce a dilation operator with $\varsigma\in\mathbb{N}$ for $\mathcal{P}$ as
@@
at time step $k$. This dialation operator robustly inflates the current parameter set $\mathcal{P}_{k}$ to account for potential changes, ensuring the true parameter at the next step $\boldsymbol{\rho}_{k+1}^{*}$ remains contained within the set. The size of this inflation is governed by the maximum parameter drift rate $\varphi_{\rho}$ from (3) making this a critical step for providing robust safety guarantees. Our robust adaptive MPC requires that we predict how the bounds $\hat{\mathcal{P}}_{\varsigma|k}$ on the time-varying parameters $\boldsymbol{\rho}_{k}$ evolve over the prediction horizon. These prediction bounds are given by
@@
for $\varsigma=1,\dots,N-1$ where $\hat{\mathcal{P}}_{\varsigma\mid k}$ is the predicted uncertainty bound at time $k+\varsigma$. Note that the recursive parameter update in (14) is designed to accommodate initial parameter uncertainty. The set-membership identification scheme actively refines the parameter estimate by iteratively tightening the admissible parameter set through newly acquired measurements. This inherent characteristic ensures that the robust safety of the system is guaranteed and performance is maintained even when commencing from a coarse initial parameter set that does not heavily rely on the precise initial set chosen.
@@
[…]
Note that Eq. (15) is an outer-approximation of Eq. (9) intended to reduce computational complexity.
@@ III-C Robust Tube Constraints
In this section, we derive a state-space tube $\mathbb{X}_{\varsigma|k}$ for $\varsigma$ and we can guarantee contains the state $\mathbf{x}_{k}$ despite parametric uncertainty $\boldsymbol{\rho}_{k}\in\mathcal{P}_{k}$ and disturbances $\mathbf{w}_{k}\in\mathbb{W}$. By remaining in the state tube, we can guarantee that the servicer would not damage the client. The states of the client and the control inputs from the servicer must satisfy the constraints as
$\mathbb{Z}$ $=\left\{(\mathbf{x}_{k},\mathbf{u}_{k})\in\mathbb{R}^{n}\times\mathbb{R}^{m}\bigr|\mathbf{F}\mathbf{x}_{k}+\mathbf{G}\mathbf{u}_{k}\leq\boldsymbol{\gamma}\right\},$ (16)
where the matrices $\mathbf{F}\in\mathbb{R}^{n_{\gamma}\times n}$ and $\mathbf{G}\in\mathbb{R}^{n_{\gamma}\times n}$ are derived from the input (4a) and state (4b) constraints. The objective is to regulate the Client state $\mathbf{x}_{k}\in\mathbb{R}^{2}$ from the initial condition $\mathbf{x}_{0}$ to the desired state $\mathbf{x}_{d}$, and the constraints the control input $\mathbf{u}_{k}\in\mathbb{R}^{6}$ from the Servicer under safety set, while robustly satisfying the constraints in (16).
The tube MPC approach proposed in [lopez2019dynamic] ensures robust constraint satisfaction. The control input is parameterized using a feedback gain $\mathbf{K}\in\mathbb{R}^{m\times n}$ as
@@
A gain $\mathbf{K}$ satisfying Assumption 2 can be computed using standard robust control techniques, e.g., following the the linear matrix inequality from [wang1992robust].
@@
for $\varsigma=0,\dots,N-1$ and for all parameters $\boldsymbol{\rho}_{k}\in\mathcal{P}_{\varsigma|k}$. This ensures that $\mathbf{x}_{\varsigma|k}\in\mathbb{X}_{\varsigma|k}$ for all the realizations of uncertainty and disturbance. To manage computational complexity, the tube cross-section at each time step, $\left\{\mathbb{X}_{\varsigma\mid k}\right\}_{j\in\{0,N\}_{\mathbb{N}}}$, is parameterized by translating and scaling of the set
@@
$\begin{aligned} \mathbb{X}_{\varsigma|k}&=\{\mathbf{z}_{\varsigma|k}\}\oplus\mathbb{X}_{0,k}^{\curlyvee}=\{\mathbf{x}\in\mathbb{R}^{n}|\mathbf{H}_{x,k}^{\curlyvee}(\mathbf{x}-\mathbf{z}_{\varsigma|k})\leq\boldsymbol{\gamma}\}\\ &=\{\mathbf{z}_{\varsigma|k}\}\oplus\vartheta_{\varsigma|k}\text{co}\{\mathbf{x}^{1},\mathbf{x}^{2},\ldots,\mathbf{x}^{v}\}.\end{aligned}$
@@ III-D Reformulation of Robust Tube Constraints
From [kouvaritakis2016model], we can reformulate the robust tube constraints (20) as linear equality and inequality constraints as follows. Let the state tube $\{\mathbb{X}_{\varsigma|k}\}_{\varsigma\in\{{0},{N}\}_{\mathbb{N}}}$ be parameterized according to (20). Then, the constraints (16) and set-dynamics (18) are satisfied if and only if $\forall j\in\{1,v\}_{\mathbb{N}}$, $\varsigma\in\{0,N{-}1\}_{\mathbb{N}}$ there exists $\boldsymbol{\Gamma}_{\varsigma|k}^{j}\in\mathbb{R}^{n_{x}\times n_{p}}_{+}$ such that
@@ III-E Robust Terminal Set
$(\mathbf{z},\vartheta)\in\mathbb{X}_{\text{T}}\Rightarrow$ $\exists\left(\mathbf{z}^{+},\vartheta^{+}\right)\in\lambda\mathbb{X}_{\text{T}}\text{ s.t. }$
@@
$(\mathbf{z},\vartheta)\in\mathbb{X}_{\text{T}}\Rightarrow$ $\exists\left(\mathbf{z}^{+},\vartheta^{+}\right)\in\mathbb{X}_{\text{T}}\text{ s.t. }$
@@
Assumption 3 implies that the set $\mathbb{X}_{T}$ is a robust positively invariant (RPI) set [kolmanovsky1998theory] for the set-dynamics in $(\mathbf{z},\vartheta)$, with an additional constraint that the set $\mathbb{X}_{N|k}$ remains centered at origin which is the desired equilibrium point $\mathbf{x}^{*}$ for dislodging problem. Note that Assumption 2 is a necessary condition for Assumption 3 to be satisfied, but they are stated separately to emphasize that the stronger assumption is only needed to implement the terminal condition.
@@ III-F Robust Adaptive MPC Algorithm
[…]
@@
Theorem 1: Suppose the Assumptions 1, 2 and 3 be satisfied and an initial feasible solution exist. Let the parameter estimate set $\mathcal{P}_{k}$ be defined by the set dynamics (13) and admissible input set $\mathcal{I}_{v_{s}}\left(\mathbf{x}_{k},\mathcal{P}_{k}\right)$ as
@@
Then, the system (6) with the proposed MPC control law in (17) using Algorithm 1 satisfies the following closed-loop properties for all $k>k_{0}$:
@@
For Property (1), we know that $\boldsymbol{\rho}_{k}\in\mathcal{P}_{k}$ and define $\epsilon_{\rho_{k}}=\boldsymbol{\rho}^{*}_{k+1}-\boldsymbol{\rho}^{*}_{k}$. Then we can get
@@
$\tilde{\mathbf{H}}_{\rho_{k}}\boldsymbol{\rho}^{*}_{k+1}=\tilde{\mathbf{H}}_{\rho_{k}}\boldsymbol{\rho}^{*}_{k}+\tilde{\mathbf{H}}_{\rho_{k}}\epsilon_{\rho_{k}}\leq\tilde{\boldsymbol{\delta}}_{\rho_{k}}+d_{\rho}\boldsymbol{\gamma}.$
Since by assumption $\boldsymbol{\rho}_{k+1}\in\mathcal{P}$, we can get $\boldsymbol{\rho}^{*}_{k+1}\in\mathbf{d}_{\varsigma}\left(\mathcal{P}_{k}\cap\Delta_{k+1}\right)$.
@@
$\mathbb{X}_{\varsigma+2\mid k}^{*}\supseteq\mathbf{A}_{\text{cl}}(\boldsymbol{\rho})\mathbb{X}_{\varsigma+1\mid k}^{*}\oplus\mathbf{B}(\boldsymbol{\rho})\mathbf{v}_{\varsigma+1\mid k}^{*}\oplus\mathbb{W}$
$\Rightarrow$ $\tilde{\mathbb{X}}_{\varsigma+1\mid k+1}\supseteq\mathbf{A}_{\text{cl}}(\boldsymbol{\rho})\tilde{\mathbb{X}}_{\varsigma\mid k+1}\oplus\mathbf{B}(\boldsymbol{\rho})\tilde{\mathbf{v}}_{\varsigma\mid k+1}\oplus\mathbb{W}.$
where $\tilde{\mathbf{v}}_{\varsigma\mid k+1}=\mathbf{v}_{\varsigma+1\mid k}^{*}$ and $\varsigma\in\{0,N{-}2\}_{\mathbb{N}}$. Based on Assumption 3 that there exist $\mathbb{X}_{N\mid k+1}$ satisfy the last set when $\varsigma=N{-}1$, that proves that $\mathcal{I}_{v_{s}}\neq\{\varnothing\}$ during all time step.
Property (3) is the direct result of Section III-D. ∎